I need help please, solving this equation 11x^2+20x-39
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

I need help please, solving this equation 11x^2+20x-39
------
It's not an equation, there's no equal sign.
It can be factored
= (11x - 13)*(x + 3)

Answer by stanbon(75887)

I need help please, solving this equation 11x^2+20x-39 = 0
-------------------------------
Factor:
(x+3)(11x-13) = 0
x = -3 or x = 13/11
=============================
Cheers,
Stan H.
Reply to stanbon@comcast.net or algebra.com